A current transformer has a single turn on the primary winding and a secondary winding of 60 turns. The secondary
winding is connected to an ammeter with a resistance of 0.15Ω. The resistance of the secondary winding is 0.25Ω. If
the current in the primary winding is 300 A, determine (a) the reading on the ammeter, (b) the potential difference
across the ammeter and (c) the total load (in VA) on the secondary.
1
Expert's answer
2020-06-18T14:50:59-0400

A) the reading on the ammeter

We know that

N1I1 = N2I2

I2 = (N1I1)/N2

= (1*300)/60

= 5 A

Hence reading of ammeter = Im = I2 = 5 A

B) the potential difference

ImRm = 5 *0.15

= 0.75 v

Hence Vm = 0.75 v

C)  the total load (in VA) on the secondary.

= V2I2

= (I2(R1+Rm))I2

= 5 (0.25+0.15) *5

= 5*0.4*5

S = 10 VA